# DPBridge: Latent Diffusion Bridge for Dense Prediction

### v3

@@ 4.5 Distribution Alignment Normalization
A key challenge during finetuning is the misalignment between the marginal distribution $q(\mathbf{z}_{t})$ of intermediate samples in DPBridge and that of the diffusion backbone, due to the difference in the forward transition kernels. This misalignment weakens the model’s denoising ability and hinders effective knowledge transfer. To address this issue, we propose a distribution alignment normalization technique that transforms intermediate latents computed from our bridge process Equation 3 to match the distributional form of a standard diffusion process.
@@
Since the square sum of coefficients is equal to 1, the normalized bridge process $\{\mathbf{z}_{t}^{\prime}:t\in[0,T]\}$ is the same as a standard VP process, making the marginal distribution of these two aligned. Therefore by feeding the normalized sample $\mathbf{z}_{t}^{\prime}$ in the network $\boldsymbol{\epsilon}_{\theta}$, we can fully exploit the denoising capability from pretrained diffusion backbone. We note that a similar strategy was independently proposed in FrameBridge [52], where it is applied in the context of image-to-video generation. Our method, developed concurrently, is specifically designed to facilitate adaptation in dense prediction tasks under the DPBridge framework. An overview of the full training and inference procedures are summarized in Algorithm 1 Algorithm 2. For more derivation details, please refer to Appendix A.
@@ 5 Experiments
We conduct evaluation primarily on two representative dense prediction tasks: depth estimation and surface normal prediction. We begin with a brief overview of experimental settings, followed by both quantitative and qualitative analyses. Next, we present ablation studies to investigate the contribution of each design component. Implementation details are presented in Appendix D for reference.
@@ 5.1 Experiments Setup
Datasets. The basic setting for depth estimation and surface normal prediction experiments are similar. We use two synthetic datasets, HyperSim [40] and Virtual KITTI [6], for training, which cover both indoor and outdoor scenes with high-quality ground-truth annotations. To comprehensively evaluate the model performance and demonstrate its generalization capability, we conduct zero-shot evaluations on real-world datasets. For depth estimation, this includes NYUv2 [44], KITTI [18], ScanNet [13], ETH3D [43] and DIODE [50]. For surface normal prediction, we additionally incorporate iBims [29] and Sintel [5] datasets. For fair comparison, we use the same validation split as in previous methods [28].
@@
DiverseDepth [58] 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1 12.2
MiDaS [38] 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5 11.8
LeReS [59] 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6 9.5
Omnidata [15] 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2 9.2
HDN [61] 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0 7.2
DPT [39] 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8 7.0
Depth Anything V2 [56] 4.4 97.9 7.5 94.8 13.2 84.2 4.4 98.0 6.5 95.4 2.3
Marigold [28] 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2 6.0
DepthFM [19] 6.5 95.6 8.3 93.4 - - - - 22.5 80.0 3.8
Geowizard [16] 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1 6.2
E2E-FT [17] 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6 3.6
I2SB [32] 23.5 63.6 35.1 46.4 53.4 40.5 21.1 67.4 55.5 58.2 14.0
BBDM [31] 33.2 49.6 58.6 25.6 52.0 40.6 25.2 61.2 70.1 46.9 15.2
DDBM [65] 26.4 62.2 55.1 27.6 69.5 33.8 23.2 65.6 52.3 56.9 14.5
@@
Table 1: Quantitative comparison on affine-invariant depth estimation benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The best and second-best results are highlighted in bold and underline formats.
@@
EESNU [3] 16.2 58.6 - - 20.0 58.5 42.1 11.5 6.8
Omnidata v1 [15] 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4 8.8
Omnidata v2 [27] 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7 6.2
DSINE [2] 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5 3.5
Marigold [28] 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9 6.4
Geowizard [16] 17.0 56.5 15.4 61.6 13.0 65.3 40.4 13.2 4.2
E2E-FT [17] 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3 2.1
StableNormal [57] 17.7 55.9 15.8 61.2 17.2 66.6 - - 5.0
@@
Table 2: Quantitative comparison on surface normal prediction benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The best and second-best results are highlighted in bold and underline formats.
@@
Figure 5: Qualitative comparison with diffusion-based surface normal prediction methods on in-the-wild images. DPBridge’s results exhibit more fine-grained geometric details, such as the curvature of architectural structures, the contour of coral and reef, and the fur texture of the cat.
@@ 5.2 Quantitative and Qualitative Results
Table 1 and Table 2 summarize the quantitative results of DPBridge on zero-shot benchmarks across depth estimation and surface normal prediction tasks. We evaluate two variants of our model—DPBridge (SD1.5) and DPBridge (SD2.1)—built on top of Stable Diffusion v1.5 and v2.1-base [41], respectively. For depth estimation, DPBridge (SD2.1) achieves an AbsRel of 4.4 and $\delta_{1}$ accuracy of 97.6% on NYUv2, and 4.5 AbsRel / 97.4% $\delta_{1}$ on ScanNet, both among the best in class. Across all five datasets, it achieves an average ranking of 2.7—second only to Depth Anything V2 (rank 2.3), which is trained on over 1,000 times more data. For surface normal prediction, DPBridge also delivers state-of-the-art performance with a mean angular error of 14.8° and $65.3\%$ accuracy under $11.25^{\circ}$ on NYUv2, outperforming both feed-forward and diffusion-based methods. It also maintains strong results on iBims-1 ($17.1$, $68.0\%$) and Sintel ($25.0$, $46.3\%$). These results highlight DPBridge’s versatility to generalize well across tasks and benchmarks. Notably, compared to prior bridge models such as I2SB, BBDM, and DDBM, DPBridge reduces average error by a large margin (e.g., 4.4 vs. 23.5 AbsRel in NYUv2 depth), underscoring the benefits and importance of incorporating pretrained visual priors to bridge construction. Furthermore, among diffusion-based approaches, DPBridge consistently achieves the best or second-best scores, validating our core hypothesis that bridge formulation provides clear advantages over standard diffusion frameworks in dense prediction tasks, and combining bridge modeling with pretrained visual prior offers a more efficient and generalizable framework.
Figure 5 and Figure 5 present qualitative comparisons of DPBridge with SOTA diffusion-based methods on in-the-wild images for depth estimation and surface normal prediction, respectively. As shown in Figure 5, DPBridge generates depth maps with sharper object boundaries and more coherent scene geometry, particularly in challenging cases such as urban scenes or complex foreground objects like animals, grass and flowers. In contrast, competing methods like Geowizard and DepthFM often produce over-smoothed outputs or fail to preserve depth discontinuities. In Figure 5, DPBridge demonstrates stronger fidelity in fine-grained geometric details, such as the curvature of architectural structures, the contour of coral and reef, and the fur texture of the cat. Competing methods like Marigold tends to focus on the center object in the image and ignore the environment, while StableNormal produces unreliable outputs. These visual results reinforce the quantitative findings and highlight the effectiveness of DPBridge in modeling structured output distributions with both global coherence and fine details.
@@ 5.3 Ablation Studies
To evaluate the effectiveness of each design component and training setting, we conduct ablation studies using depth estimation tasks on the NYUv2 and KITTI datasets. The results are summarized in Table 3 and Figure 6, with detailed discussions provided below.
Visual Prior Exploitation. To evaluate the usage of visual priors, we compare finetuning with training from scratch. As shown in the first row of Table 3, finetuning yields significantly better results under identical settings. This is attributed to stronger visual priors from foundation models and faster convergence. Additionally, the last rows of Table 1 and Table 2 show that SD2.1 outperforms SD1.5, highlighting that higher-quality priors further boost performance. These results confirm the advantage of building DPBridge on strong pretrained backbones.
Distribution Alignment Normalization. We next assess the benefit of aligning marginal distributions between the diffusion bridge and standard diffusion processes. As shown in the second and fourth rows of Table 3, adding a distribution alignment normalization step before the noise prediction network consistently improves performance.
Image Consistency Loss. As shown in the third and fourth rows of Table 3, the auxiliary image consistency loss improves performance but has a smaller impact than distribution alignment. This is likely because it mainly affects the boundaries or overlapping regions between masked and valid areas, which are limited in scope. Nonetheless, when combined with other components, it contributes to a notable overall performance gain.
Number of Sampling Steps. We evaluate performance under varying sampling steps during inference. As shown in Figure 6, perception quality improves rapidly with more steps and saturates after 10. Notably, compared to Table 1, where other diffusion-based methods use 50 steps, DPBridge achieves comparable results with significantly fewer steps, highlighting its efficiency.
@@ Appendix A Derivation Details
In this section, we’ll provide derivation details of the maximum likelihood loss presented in Section 4.4. Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Equation 5, whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},z\mathbf{z}_{T})$. The last two term of the posterior can be retrieved directly through the marginal distribution of diffusion bridge’s forward process, but further derivation is required for the first term. Specifically, as specified in Equation 3:
@@
To derive the transition probability $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, we substitute $\mathbf{z}_{0}$ in Equation 12 with derivation from Equation 13, obtain
@@
Putting all these terms back to Equation 5, we’ll have:
@@
$\propto$ $\exp\bigg{[}-\frac{1}{2}\bigg{(}\frac{(\mathbf{z}_{t}-a_{t}\mathbf{z}_{t-1}-b_{t}\mathbf{z}_{T})^{2}}{\delta_{t}^{2}}+\frac{(\mathbf{z}_{t-1}-m_{t-1}\mathbf{z}_{0}-n_{t-1}\mathbf{z}_{T}^{2}}{\sigma_{t-1}^{2}}-\frac{(\mathbf{z}_{t}-m_{t}\mathbf{z}_{0}-n_{t}\mathbf{z}_{T}^{2}}{\sigma_{t}^{2}}\bigg{)}\bigg{]}$
@@
whose results are equivalent to Equation 6 and Equation 7.
@@
which is the same as in Equation 8. This concludes our derivation.
@@ B.2 Accelerate Inference
Similar to DDIM [45] and DBIM [64], the inference process of DPBridge can be accelerated using non-Markovian sampling scheme that preserves the marginal distributions of intermediate states. Specifically, we perform sampling over a reduced set of timesteps $\{t_{i}\}_{i=1}^{N}$, where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$, allowing the number of inference steps $N$ to be decoupled from the total number of training steps $T$. The accelerated inference process is given by:
@@
where $g_{n}$ denotes variance parameter that governs the stochasticity of the reverse sampling process. Notably, when $g_{n}=\tilde{\sigma}_{t}^{2}$, as defined Equation 5, the process becomes a Markovian bridge.

### v4

@@ 4.5 Distribution Alignment Normalization
A key challenge during finetuning is the misalignment between the marginal distribution $q(\mathbf{z}_{t})$ of intermediate samples in DPBridge and that of the diffusion backbone, due to the difference in the forward transition kernels. This misalignment weakens the model’s denoising ability and hinders effective knowledge transfer. To address this issue, we propose a distribution alignment normalization technique that transforms intermediate latents computed from our bridge process Equation˜3 to match the distributional form of a standard diffusion process.
@@
Since the square sum of coefficients is equal to 1, the marginal distribution of the normalized bridge process $\{\mathbf{z}_{t}^{\prime}:t\in[0,T]\}$ can thus align with that of the standard diffusion process, while preserving bridge-specific dynamics. Therefore, by feeding the normalized sample $\mathbf{z}_{t}^{\prime}$ in the network $\boldsymbol{\epsilon}_{\theta}$, we can fully exploit the denoising capability from the pretrained diffusion backbone. We note that a similar strategy was independently proposed in FrameBridge [wang2024framebridge], where it is applied in the context of image-to-video generation. Our method, developed concurrently, is specifically designed to facilitate adaptation in dense prediction tasks under the DPBridge framework. An overview of the full training and inference procedures is summarized in Algorithm˜1 Algorithm˜2.
@@ 5 Experiments
We conduct evaluation primarily on two representative dense prediction tasks: depth estimation and surface normal prediction. We begin with a brief overview of experimental settings, followed by both quantitative and qualitative analyses. Next, we present ablation studies to investigate the contribution of each design component. Implementation details are presented in Appendix˜D for reference.
@@ 5.1 Experiments Setup
Datasets. The basic setting for depth estimation and surface normal prediction experiments are similar. We use two synthetic datasets, HyperSim [roberts2021hypersim] and Virtual KITTI [cabon2020virtual], for training, which cover both indoor and outdoor scenes with high-quality ground-truth annotations. To comprehensively evaluate the model performance and demonstrate its generalization capability, we conduct zero-shot evaluations on real-world datasets. For depth estimation, this includes NYUv2 [silberman2012indoor], KITTI [geiger2012we], ScanNet [dai2017scannet], ETH3D [schops2017multi] and DIODE [vasiljevic2019diode]. For surface normal prediction, we additionally incorporate iBims [koch2018evaluation] and Sintel [butler2012naturalistic] datasets. For fair comparison, we use the same validation split as in previous methods [ke2024repurposing].
@@
DiverseDepth [yin2020diversedepth] 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1 12.2
MiDaS [ranftl2020towards] 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5 11.8
LeReS [yin2021learning] 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6 9.5
Omnidata [eftekhar2021omnidata] 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2 9.2
HDN [zhang2022hierarchical] 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0 7.2
DPT [ranftl2021vision] 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8 7.0
Depth Anything V2 [yang2024depth] 4.4 97.9 7.5 94.8 13.2 84.2 4.4 98.0 6.5 95.4 2.3
Marigold [ke2024repurposing] 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2 6.0
DepthFM [gui2024depthfm] 6.5 95.6 8.3 93.4 - - - - 22.5 80.0 3.8
Geowizard [fu2025geowizard] 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1 6.2
E2E-FT [garcia2024fine] 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6 3.6
I2SB [liu20232] 23.5 63.6 35.1 46.4 53.4 40.5 21.1 67.4 55.5 58.2 14.0
BBDM [li2023bbdm] 33.2 49.6 58.6 25.6 52.0 40.6 25.2 61.2 70.1 46.9 15.2
DDBM [zhou2023denoising] 26.4 62.2 55.1 27.6 69.5 33.8 23.2 65.6 52.3 56.9 14.5
@@
Table 1: Quantitative comparison on affine-invariant depth estimation benchmarks. All metrics are presented in percentage terms. $F$, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The best and second-best results are highlighted in bold and underline formats.
@@
EESNU [bae2021estimating] 16.2 58.6 - - 20.0 58.5 42.1 11.5 6.8
Omnidata v1 [eftekhar2021omnidata] 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4 8.8
Omnidata v2 [kar20223d] 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7 6.2
DSINE [bae2024rethinking] 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5 3.5
Marigold [ke2024repurposing] 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9 6.4
Geowizard [fu2025geowizard] 17.0 56.5 15.4 61.6 13.0 65.3 40.4 13.2 4.2
E2E-FT [garcia2024fine] 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3 2.1
StableNormal [ye2024stablenormal] 17.7 55.9 15.8 61.2 17.2 66.6 - - 5.0
@@
Table 2: Quantitative comparison on surface normal prediction benchmarks. All metrics are presented in percentage terms. $F$, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The best and second-best results are highlighted in bold and underline formats.
@@
Figure 5: Qualitative comparison with diffusion-based surface normal prediction methods on in-the-wild images. DPBridge exhibits more fine-grained geometric details, such as the coral and reef’s contour, the kitten’s fur details, and the robot’s body composition curvature.
@@ 5.2 Quantitative and Qualitative Results
Tab.˜1 and Tab.˜2 summarize the quantitative results of DPBridge on zero-shot benchmarks across depth estimation and surface normal prediction tasks. We evaluate two variants of our model—DPBridge (SD1.5) and DPBridge (SD2.1)—built on top of Stable Diffusion v1.5 and v2.1-base [rombach2022high], respectively. For depth estimation, DPBridge (SD2.1) achieves an AbsRel of 4.4 and $\delta_{1}$ accuracy of 97.6% on NYUv2, and 4.5 AbsRel / 97.4% $\delta_{1}$ on ScanNet, both among the best in class. Across all five datasets, it achieves an average ranking of 2.7—second only to Depth Anything V2 (rank 2.3), which is trained on over 1,000 times more data. For surface normal prediction, DPBridge also delivers state-of-the-art performance with a mean angular error of 14.8° and $65.3\%$ accuracy under $11.25^{\circ}$ on NYUv2, outperforming both feed-forward and diffusion-based methods. It also maintains strong results on iBims-1 ($17.1$, $68.0\%$) and Sintel ($25.0$, $46.3\%$). These results highlight DPBridge’s versatility to generalize well across tasks and benchmarks. Notably, compared to prior bridge models such as I2SB, BBDM, and DDBM, DPBridge reduces average error by a large margin (e.g., 4.4 vs. 23.5 AbsRel in NYUv2 depth), underscoring the benefits and importance of incorporating pretrained visual priors to bridge construction. Furthermore, among diffusion-based approaches, DPBridge consistently achieves the best or second-best scores, validating our core hypothesis that bridge formulation provides clear advantages over standard diffusion frameworks in dense prediction tasks, and combining bridge modeling with pretrained visual prior offers a more efficient and generalizable framework.
Figure˜5 and Figure˜5 present qualitative comparisons of DPBridge with SOTA diffusion-based methods on in-the-wild images for depth estimation and surface normal prediction, respectively. As shown in Figure˜5, DPBridge generates depth maps with sharper object boundaries and more coherent scene geometry, particularly in challenging cases such as urban scenes or complex foreground objects like animals, grass, and flowers. In contrast, competing methods like Geowizard and DepthFM often produce over-smoothed outputs or fail to preserve depth discontinuities. In Figure˜5, DPBridge demonstrates stronger fidelity in fine-grained geometric details, such as the contour of coral and reef, the fur details of the cat, and the body composition curvature of the robot. Competing method like Marigold tends to focus on the center object in the image and ignore the environment, while StableNormal produces unreliable outputs. These visual results reinforce the quantitative findings and highlight the effectiveness of DPBridge in modeling structured output distributions with both global coherence and fine details.
@@ 5.3 Ablation Studies
To evaluate the effectiveness of each design component and training setting, we conduct ablation studies using depth estimation tasks on the NYUv2 and KITTI datasets. The results are summarized in Tab.˜3 and Figure˜6, with detailed discussions provided below.
Visual Prior Exploitation. To evaluate the usage of visual priors, we compare finetuning with training from scratch. As shown in the first row of Tab.˜3, finetuning yields significantly better results under identical settings. This is attributed to stronger visual priors from foundation models and faster convergence. Additionally, the last rows of Tab.˜1 and Tab.˜2 show that SD2.1 outperforms SD1.5, highlighting that higher-quality priors further boost performance. These results confirm the advantage of building DPBridge on strong pretrained backbones.
Distribution Alignment Normalization. We next assess the benefit of aligning marginal distributions between the diffusion bridge and standard diffusion processes. As shown in the second and fourth rows of Tab.˜3, adding a distribution alignment normalization step before the noise prediction network consistently improves performance.
Image Consistency Loss. As shown in the third and fourth rows of Tab.˜3, the auxiliary image consistency loss improves performance but has a smaller impact than distribution alignment. This is likely because it mainly affects the boundaries or overlapping regions between masked and valid areas, which are limited in scope. Nonetheless, when combined with other components, it contributes to a notable overall performance gain.
Number of Sampling Steps. We evaluate performance under varying sampling steps during inference. As shown in Figure˜6, perception quality improves rapidly with more steps and saturates after 10. Notably, compared to Tab.˜1, where other diffusion-based methods use 50 steps, DPBridge achieves comparable results with significantly fewer steps, highlighting its efficiency.
@@ 6 Conclusion
Acknowledgement
This research is funded in part by an ARC (Australian Research Council) Discovery Grant of DP220100800. Prof Hongdong Li holds concurrent appointments as a Full Professor with the ANU and as an Amazon Scholar with Amazon (parttime). This paper describes work performed at ANU and is not associated with Amazon.
@@ Appendix A Derivation Details
In this section, we’ll provide derivation details of the maximum likelihood loss presented in Sec.˜4.4. Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Equation˜5, whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})$. The last two terms of the posterior can be retrieved directly through the marginal distribution of the diffusion bridge’s forward process, but further derivation is required for the first term. Specifically, as specified in Equation˜3:
@@
To derive the transition probability $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, we substitute $\mathbf{z}_{0}$ in Equation˜12 with derivation from Equation˜13, obtain
@@
Putting all these terms back to Equation˜5, we’ll have:
@@
$\propto$ $\exp\bigg[-\frac{1}{2}\bigg(\frac{(\mathbf{z}_{t}-a_{t}\mathbf{z}_{t-1}-b_{t}\mathbf{z}_{T})^{2}}{\delta_{t}^{2}}+\frac{(\mathbf{z}_{t-1}-m_{t-1}\mathbf{z}_{0}-n_{t-1}\mathbf{z}_{T})^{2}}{\sigma_{t-1}^{2}}-\frac{(\mathbf{z}_{t}-m_{t}\mathbf{z}_{0}-n_{t}\mathbf{z}_{T})^{2}}{\sigma_{t}^{2}}\bigg)\bigg]$
@@
whose results are equivalent to Equation˜6 and Equation˜7.
@@
which is the same as in Equation˜8. This concludes our derivation.
@@ B.2 Accelerate Inference
Similar to DDIM [song2020denoising] and DBIM [zheng2024diffusion], the inference process of DPBridge can be accelerated using non-Markovian sampling scheme that preserves the marginal distributions of intermediate states. Specifically, we perform sampling over a reduced set of timesteps $\{t_{i}\}_{i=1}^{N}$, where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$, allowing the number of inference steps $N$ to be decoupled from the total number of training steps $T$. The accelerated inference process is given by:
@@
where $g_{n}$ denotes variance parameter that governs the stochasticity of the reverse sampling process. Notably, when $g_{n}=\tilde{\sigma}_{t}^{2}$, as defined Equation˜5, the process becomes a Markovian bridge.